OpenStudy (anonymous):

The graph of which quadratic equation has a vertex of (–2, –1)?

OpenStudy (anonymous):

Choices: y = (x - 2)^2 - 1 y = (x + 2)^2 - 1 y = (x - 1)^2 - 2 y = (x + 1)^2 - 2

terenzreignz (terenzreignz):

Doesn't get much simpler than this :D A quadratic equation of the form y = (x - h)² + k has its vertex at the point (h,k)

OpenStudy (anonymous):

haha okay

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

i got b

terenzreignz (terenzreignz):

And you're right :)